OpenStudy (anonymous):

Solve the following for θ where θ is greater than or equal to 0 and less than 2π: cos^2θ=3cosθ

OpenStudy (dumbcow):

move everything to left side so it equals 0 factor out a cos set each factor equal to 0 and solve

OpenStudy (anonymous):

Would the answer for \[\theta \] be 0 then?

OpenStudy (anonymous):

So if cosθ=0, it's just π/2 and 3π/2?

OpenStudy (jdoe0001):

\(\bf cos(2 \theta) = 3cos(\theta) \implies cos(2 \theta)-3cos(\theta) = 0\\ \color{blue}{cos(2 \theta) = 2cos^2(\theta) -1}\\ cos(2 \theta)-3cos(\theta) = 0 \implies 2cos^2(\theta) -1-3cos(\theta) = 0\\ 2cos^2(\theta)-3cos(\theta)-1=0 \implies 2x^2-3x-1=0\) so solve it, I gather you may end up using the quadratic formula for that

OpenStudy (dumbcow):

yes its pi/2 and 3pi/2 im assuming its cos squared correct ?